# Why E.T. Can’t Phone Home: A Global View on IP-based Geoblocking at VoWiFi

### v1

@@ 4.2.1. Measurement Coverage and Domestic Results
Our measurements are limited by the countries that are available via our VPN- and cloud services. We cover 107/109 (IPv4) and 16/16 (IPv6) of our target countries (target territories defined by the DNS discovery). The two countries we are missing for IPv4 are both overseas departments of France, which are relatively small countries: French Polynesia (one operator) and La Réunion (two operators). Within French Polynesia, the single operator that was discovered via DNS (Ora${}_{[54705]}$) responds to IKE requests from all over the globe and thereby is not geoblocked. The same holds true for one Reunionese carrier(Zeop${}_{[64704]}$), while the second one (Orange${}_{[64700]}$) was not responsive from any country we scanned from and thereby is out of scope of our measurement coverage [^9] For the remaining paper, we’ve treated it as a non-responsive ePDG, although technically it could be geoblocked and merely reachable via the home country (La Reunion).. Figure 5 gives an overview of the global scope of our measurements: We’ve accomplished dual-stack connectivity in nearly all covered countries (1) and have some countries with IPv4 only coverage (2). Several countries that were not available via our scan infrastructure do not have VoWiFi yet (3), i.e., there are no DNS entries for any ePDG within that area. The two countries with missing coverage (i.e., French Polynesia and La Réunion) are very small and thereby not visible on the global map (4).
@@
Region (via MCC) Countries ${}^{\mathrm{b}}$ Domains IPs Countries ${}^{\mathrm{b}}$ Domains IPs
@@ 9 Worldwide ${}^{\mathrm{c}}$ 0 0% 0 0% 0 0% 0 - 0 - 0 -
${}^{\mathrm{b}}$ according to the MCC.   ${}^{\mathrm{c}}$ Satellite, Air, Maritime, Antarctica.
@@
Dealing with Inaccuracies. Besides blocking according to a user’s location, operators might also block the usage of popular VPN services or cloud infrastructure as a connection relay. More specifically, they might block some of our VPN services, or Autonomous Systems (ASes) that are commonly shared among popular VPN services. Furthermore, not all operators use MaxMind as their geolocation broker, introducing potential inconsistencies between our classification and the actual geolocation at the operator’s sides. Additionally, we might occasionally experience connection hiccups within our scan infrastructure (e.g., when the used VPN unexpectedly closes the current connection). Lastly, some ePDGs seem to be very sensitive and do not accept repeated connections from the same IP address, simply ignore IKE handshakes that contain inappropriate cipher sets or go on/offline (i.e., experience downtimes) throughout our measurement study. To cope with these instabilities, we randomly switch VPN servers between subsequent measurements and aim for a big and diverse set of VPN services and realized measurements. Also, we use the responsiveness (i.e., the percentage an ePDG was reported as responsive) of the domestic case as a baseline, to decide whether the server was blocked in a roaming case. As a matter of precaution, we consider some space for generous error bounds (e.g., when we or the operator misclassify the VPN IP and assume the wrong country) by only flagging an area as geoblocked, when it has less than 10% of the assumed baseline responsiveness. Thereby, the responsiveness remains comparable and we can correctly classify an ePDG, even when unexepcted events occur, e.g., when the ePDG server experiences a downtime during our measurement campaign.
@@ 4.2.2. Roaming Results
Interestingly, we found numerous operators in European countries (Austria, Czech Republic, Denmark, France, Hungary, Luxembourg, the Netherlands, Norway, Romania, Sweden and the United Kingdom) with large-scale blocking measures. Moreover, many operators within the EU also block connections from their neighbouring EU-countries. In contrast, intra-EU roaming via regular radio access is possible without additional costs in these countries, due to the Roam Like At Home (RLAH) doctrine (EU-2022-612, 15). As an exception, we’ve found a single Slovakian operator with worldwide geoblocking that specifically exempts EU/EAA countries from the blocking.
@@
Region (via MCC) Countries ${}^{\mathrm{b}}$ Domains Countries ${}^{\mathrm{b}}$ Domains
@@ 7 South- & Central America 0 0% 0 0% 0 0% 0 0%
Interestingly, we’ve found one Hungarian operator that supports both IPv4 and IPv6 and blocks external connections only on the IPv4 stack. Thereby, customers could circumvent the blocking by simply connecting via IPv6 (a similar phenomenon has been discovered by previous research  (czyz2016backdoor, 8)).
@@
Revisited: Reliance Jio${}_{[405874]}$ India. As stated in Section  G3, we found that this operator clearly separates the IP addresses that are returned via DNS for local and external customers. Analyzing the two subsets (i.e., local, and external) at the ePDG layer, we see that this behavior also reoccurs when connecting to the gateway: The set of local IP addresses does not accept any connections from abroad. Interestingly, this behavior is also mutual, i.e., the external ePDG does not accept any local connections from India either. Since a customer can thereby establish a connection to an ePDG from any location we did not account this behaviour as geoblocking. However, the existing separation mechanism could possibly be used to differentiate or block at a later stage.
Revisited: Vodafone${}_{[26202]}$ Germany. In Section  G4 we showed that this operator uses DNS-based blocking to prevent customers from connecting to the ePDG from external locations. Taking a closer look at the probing results of the corresponding endpoints, we see that the blocking is solely done by the DNS server and does not occur at the IKE layer (i.e., it accepts connections from all tested countries). Thereby, sneaky customers may simply evade the blocking by manually adding the correct DNS entries to their system, or by using specific DNS servers that always return the corresponding IPs (e.g., a local resolver in Germany that does not forward the client’s subnet).
@@ 5. Related Work
Geoblocking and Internet Censorship. In 2018, the European Union (EU) banned unjustified geoblocking within the European single market (EU-2018-302, 14). Nevertheless, the regulation includes a number of exceptions (e.g., for copyrighted audiovisual content like Netflix) and significant portions of the world remain without regulatory measures.
@@
McDonald et al. (McDonald2018Forbidden, 38) proved the prevalence of geoblocking practices in the Internet by finding geoblocking at large CDNs for nearly all of the 177 examined countries.
@@
Additionally, Kumar et al. (Kumar2022Large, 29) analyzed the mobile app ecosystem from vantage points in 26 countries. Aside from geoblocking being a common practice in the mobile app field they also found geodifferences occurring between differing countries (i.e., developers shipping different versions of an app to specific regions).
@@
Lastly, geoblocking is often introduced as a governmental censorship measure. Ramesh et al. (Ramesh2023Network, 45) showed, that — ever since Russia’s invasion of Ukraine in February 2022 — there are geofences between Russia and the rest of the world. Thereby, Russian users are not able to consume Western news or social media and Russian government domains remain inaccessible from regions like the EU and US.
@@
VoLTE and VoWiFi Security. Prior research discovered numerous security- and privacy-related vulnerabilities in VoLTE and VoWiFi. For example, Kim (Kim2015Breaking, 27) showed that early VoLTE deployments were prone to data traffic free-riding attacks since the packet-switched voice channel provided an unmetered breakout to the public Internet. Furthermore, both VoLTE and VoWiFi were found to occasionally leak precise subscriber info (e.g., Cell IDs) via the underlying SIP traffic (Kim2015Tracking, 28). Additionally, VoWiFi is vulnerable to IMSI catching attacks (OHanlon2016Blackhat, 40, 41).
@@
More recently, Lu et al. (Lu2020Ghost, 31), Xie et al. (Xie2020Untold, 54), and Lee et al. (Lee2022VWAnalyzer, 30) presented practical denial of service (DoS) attacks for VoLTE and/or VoWiFi. Moreover, Hu et al. showed that VoLTE’s emergency services are also vulnerable to DoS and free-riding attacks (Hu2022Uncovering, 20).
@@
In 2023, the Google Project Zero team discovered four severe Exynos vulnerabilities that allowed an attacker to execute arbitrary commands on the baseband processor of the most recent Pixel and many Samsung phones by injecting malicious SIP messages (Google2023Exynos, 39) into the VoLTE/VoWiFi VoIP traffic.
@@
Active Roaming Experiments. Large-scale studies that involve many operators and countries are usually limited by the complex ecosystem and the required coordination effort. However, there are several approaches (Mandalari2018Experience, 35, 49) where specifically built measurement devices were placed into target locations to measure the implications (e.g., QoE) of roaming. Additionally, Sahin and Francillon (Sahin2016Over, 48) observed hijacking of traditional voice calls that were redirected to over-the-top (OTT) services (e.g., WhatsApp, Viber) to bypass/monetize termination fees.
@@
Recently, Gegenhuber et al. (Gegenhuber2023MobileAtlas, 18, 17) introduced the MobileAtlas measurement platform that tries to overcome the mentioned scalability issues by tunneling the communication between SIM card and modem over the Internet. Their platform provides flexible roaming measurements and capabilities for a rich set of cellular features, including Internet and voice-based measurements.
@@
Evaluating and Fingerprinting VPNs. The commercial VPN ecosystem is a multi-billion dollar industry (VPNDemand, 19) and thereby has been an interesting research target. For example, previous work (Khan2018Empirical, 26, 44) analyzed and compared existing VPN services and found that many solutions leak user traffic or advertise wrong server locations.
@@
In contrast, Maghsoudlou et al. (Maghsoudlou2023Characterizing, 34) did not purchase any VPN subscriptions but executed Internet-wide scans to discover and fingerprint VPNs, finding over 7 million IPsec servers.
@@ 6.1. Implications to Emergency Calling
Recent reports show, that there currently is no adequate support for VoLTE roaming in substantial parts of the world (Capacitymedia2022VoLTE1, 46, 52). Since many operators are actively shutting down their 2G/3G legacy networks, this scenario has the potential to result in significant repercussions for the functionality of emergency calling (Hendrik2022VoLTE, 47, 10, 11).
@@
Although we believe that affected operators will address and fix these issues in the long term, VoWiFi could help to mitigate these shortcomings in the present day. Tourists or travellers frequently have access to WiFi, such as in their accommodations or through free WiFi hotspots in public spaces. According to the specification (EtsiArchitectureNon3gpp, 13), VoWiFi should be used by the UE for emergency calling when traditional radio services (VoLTE roaming, CSFB) are not available. The phone ultimately tries to reach both the home and the visited ePDG. However, there are circumstances where the visited ePDG might not be available, e.g., when there is no VoWiFi support in the visited country or when the customer leaves the phone in flight mode to not cause any unintentional roaming fees and thereby the current location is not known to the phone).
@@
While operators can override the default ePDG for emergency services by setting corresponding DNS records (sos.epdg.epc.mnc{y}.mcc{x}.pub.3gppnetwork.org) (EtsiNumberingAdressingIdentification, 12), only four operators had appropriate DNS entries. In all four cases, the emergency ePDG was referencing the IP address of the original ePDG, which is also the default behaviour when no sos entry is found.
@@
Thus, if an operator deploys DNS- or IKE-based geoblocking, these measures will also impact emergency services, actively denying persons in need from making an emergency call. Similarly, the emergency service via Wi-Fi is also unavailable in the home country, when customers use a VPN connection or an international SIM card that provides the Internet uplink via a foreign country.
@@ 6.2. Limitations
Although we are able to detect a considerable amount of blocking operators, there are several factors that limit our approach. Due to the fact that these constraints limit the geoblocking variants that we’re able to detect, our results can be seen as a lower bound. Thereby, the actual occurrence of geoblocking in practice is most likely even higher.
@@
More Advanced Blocking. Our current approach is designed to detect simple IP-based blocking rules. However, some operators use more sophisticated (e.g., IMEI-based) ways to decide whether a customer should be able to use VoWiFi under current circumstances. For example, an operator could use the last known roaming status via the radio network to decide whether a customer is currently in their home country or abroad. Additionally, VoWiFi roaming can only be offered to a limited set of customers (i.e., premium users). For example, we found an Austrian operator that requires an additional subscription package to get VoWiFi enabled during international roaming (Magenta2023VoWiFiRoaming, 33). In this case, the operator initially accepts IKE packets from all locations and decides whether to grant (or drop) access to VoWiFi at a later stage, when the subscriber’s identity is known. We focus on straightforward IP-based blocking, because measuring these more advanced blocking techniques on a global scale would necessitate the acquisition of SIM cards for an unfeasible number of operators. Moreover, it is crucial to highlight that relying solely on IP address-based blocking for VoWiFi access also results in the restriction of access to emergency calling services. Implementing blocking techniques that compromise the unconditional availability of emergency services that people depend on in life-threatening situations is highly ill-advised.
@@
Implementation Incompatibility & Blocked VPN Services. Our IKE probing is based on a cellular-specific open-source implementation of the IKEv2 protocol [^10] https://github.com/fasferraz/SWu-IKEv2. Possibly, some ePDGs are not compatible with this implementation or do not answer requests that offer inappropriate cipher suites. Similarly, operators could block access from well-known VPN services or ASes and IP ranges that are used by them. As a countermeasure, we use multiple services to increase the diversity within our clients. Since we got responses from a high share of ePDGs that were discovered via DNS, these two factors do not considerably influence our results.
@@
Non-Native and Non-3GPP-Compliant VoWiFi. Whereas our approach focuses on the native 3GPP version of VoWiFi there are also other ways to support voice calling functionalities over Wi-Fi networks. For example, an operator might require their customers to download and install a dedicated VoWiFi app that gives direct access to the IMS network in a non-compliant way (GSMA2017ChinaVoWiFi, 32). Additionally, some operators might use non-default ways to communicate their ePDGs, e.g., only resolve the hostname via an internal DNS server that is shipped to their customers via DHCP. Lastly, we do not cover VoIP calling via OTT services (e.g., WhatsApp, Viber, Skype).

### v2

@@ 4.2.1. Measurement Coverage and Domestic Results
Our measurements are limited by the countries that are available via our VPN- and cloud services. We cover 107/109 (IPv4) and 16/16 (IPv6) of our target countries (target territories defined by the DNS discovery). The two countries we are missing for IPv4 are both overseas departments of France, which are relatively small countries: French Polynesia (one operator) and La Réunion (two operators). Within French Polynesia, the single operator that was discovered via DNS (Ora[54705]) responds to IKE requests from all over the globe and thereby is not geoblocked. The same holds true for one Reunionese carrier(Zeop[64704]), while the second one (Orange[64700]) was not responsive from any country we scanned from and thereby is out of scope of our measurement coverage [^9] For the remaining paper, we’ve treated it as a non-responsive ePDG, although technically it could be geoblocked and merely reachable via the home country (La Reunion).. Figure 5 gives an overview of the global scope of our measurements: We’ve accomplished dual-stack connectivity in nearly all covered countries (1) and have some countries with IPv4 only coverage (2). Several countries that were not available via our scan infrastructure do not have VoWiFi yet (3), i.e., there are no DNS entries for any ePDG within that area. The two countries with missing coverage (i.e., French Polynesia and La Réunion) are very small and thereby not visible on the global map (4).
@@
Region (via MCC) Countries b Domains IPs Countries b Domains IPs
@@ 9 Worldwide c 0 0% 0 0% 0 0% 0 - 0 - 0 -
b according to the MCC.   c Satellite, Air, Maritime, Antarctica.
@@
Dealing with Inaccuracies. Besides blocking according to a user’s location, operators might also block the usage of popular VPN services or cloud infrastructure as a connection relay. More specifically, they might block some of our VPN services, or Autonomous Systems (ASes) that are commonly shared among popular VPN services. Furthermore, not all operators use MaxMind as their geolocation broker, introducing potential inconsistencies between our classification and the actual geolocation at the operator’s sides. Additionally, we might occasionally experience connection hiccups within our scan infrastructure (e.g., when the used VPN unexpectedly closes the current connection). Lastly, some ePDGs seem to be very sensitive and do not accept repeated connections from the same IP address, simply ignore IKE handshakes that contain inappropriate cipher sets or go on/offline (i.e., experience downtimes) throughout our measurement study. To cope with these instabilities, we randomly switch VPN servers between subsequent measurements and aim for a big and diverse set of VPN services and realized measurements.
Also, we use the responsiveness (i.e., the percentage an ePDG was reported as responsive) of the domestic case as a baseline, to decide whether the server was blocked in a roaming case. As a matter of precaution, we consider some space for generous error bounds (e.g., when we or the operator misclassify the VPN IP and assume the wrong country) by only flagging an area as geoblocked, when it has less than 10% of the assumed baseline responsiveness. Thereby, the responsiveness remains comparable and we can correctly classify an ePDG, even when unexepcted events occur, e.g., when the ePDG server experiences a downtime during our measurement campaign. Also, this approach allows us to more accurately flag a country as geoblocked, even when we get responses in a minority of cases due to geolocation errors from the operator.
@@ 4.2.2. Roaming Results
Interestingly, we found numerous operators in European countries (Austria, Czech Republic, Denmark, France, Hungary, Luxembourg, the Netherlands, Norway, Romania, Sweden, and the United Kingdom) with large-scale blocking measures. Moreover, many operators within the EU also block connections from their neighboring EU countries. In contrast, intra-EU roaming via regular radio access is possible without additional costs in these countries, due to the Roam Like At Home (RLAH) doctrine (EU-2022-612). As an exception, we’ve found a single Slovakian operator with worldwide geoblocking that specifically exempts EU/EAA countries from the blocking.
@@
Region (via MCC) Countries b Domains Countries b Domains
@@ 7 South- & Central America 0 0% 0 0% 0 0% 0 0%
Interestingly, we’ve found one Hungarian operator that supports both IPv4 and IPv6 and blocks external connections only on the IPv4 stack. Thereby, customers could circumvent the blocking by simply connecting via IPv6 (a similar phenomenon has been discovered by previous research  (czyz2016backdoor)).
@@
Revisited: Reliance Jio[405874] India. As stated in Section  G3, we found that this operator clearly separates the IP addresses that are returned via DNS for local and external customers. Analyzing the two subsets (i.e., local, and external) at the ePDG layer, we see that this behavior also reoccurs when connecting to the gateway: The set of local IP addresses does not accept any connections from abroad. Interestingly, this behavior is also mutual, i.e., the external ePDG does not accept any local connections from India either. Since a customer can thereby establish a connection to an ePDG from any location we did not account this behaviour as geoblocking. However, the existing separation mechanism could possibly be used to differentiate or block at a later stage.
Revisited: Vodafone[26202] Germany. In Section  G4 we showed that this operator uses DNS-based blocking to prevent customers from connecting to the ePDG from external locations. Taking a closer look at the probing results of the corresponding endpoints, we see that the blocking is solely done by the DNS server and does not occur at the IKE layer (i.e., it accepts connections from all tested countries). Thereby, sneaky customers may simply evade the blocking by manually adding the correct DNS entries to their system, or by using specific DNS servers that always return the corresponding IPs (e.g., a local resolver in Germany that does not forward the client’s subnet).
@@ 5. Related Work
Geoblocking and Internet Censorship. In 2018, the European Union (EU) banned unjustified geoblocking within the European single market (EU-2018-302). Nevertheless, the regulation includes a number of exceptions (e.g., for copyrighted audiovisual content like Netflix) and significant portions of the world remain without regulatory measures.
@@
McDonald et al. (McDonald2018Forbidden) proved the prevalence of geoblocking practices in the Internet by finding geoblocking at large CDNs for nearly all of the 177 examined countries.
@@
Additionally, Kumar et al. (Kumar2022Large) analyzed the mobile app ecosystem from vantage points in 26 countries. Aside from geoblocking being a common practice in the mobile app field they also found geodifferences occurring between differing countries (i.e., developers shipping different versions of an app to specific regions).
@@
Lastly, geoblocking is often introduced as a governmental censorship measure. Ramesh et al. (Ramesh2023Network) showed, that — ever since Russia’s invasion of Ukraine in February 2022 — there are geofences between Russia and the rest of the world. Thereby, Russian users are not able to consume Western news or social media and Russian government domains remain inaccessible from regions like the EU and US.
@@
VoLTE and VoWiFi Security. Prior research discovered numerous security- and privacy-related vulnerabilities in VoLTE and VoWiFi. For example, Kim (Kim2015Breaking) showed that early VoLTE deployments were prone to data traffic free-riding attacks since the packet-switched voice channel provided an unmetered breakout to the public Internet. Furthermore, both VoLTE and VoWiFi were found to occasionally leak precise subscriber info (e.g., Cell IDs) via the underlying SIP traffic (Kim2015Tracking). Additionally, VoWiFi is vulnerable to IMSI catching attacks (OHanlon2016Blackhat; OHanlon2017Mobile).
@@
More recently, Lu et al. (Lu2020Ghost), Xie et al. (Xie2020Untold), and Lee et al. (Lee2022VWAnalyzer) presented practical Denial of Service (DoS) attacks for VoLTE and/or VoWiFi. Moreover, Hu et al. showed that VoLTE’s emergency services are also vulnerable to DoS and free-riding attacks (Hu2022Uncovering).
@@
In 2023, the Google Project Zero team discovered four severe Exynos vulnerabilities that allowed an attacker to execute arbitrary commands on the baseband processor of the most recent Pixel and many Samsung phones by injecting malicious SIP messages (Google2023Exynos) into the VoLTE/VoWiFi VoIP traffic.
Lastly, Gegenhuber et al. (Gegenhuber2024Never; Gegenhuber2024Diffie) uncovered insecure VoWiFi configurations and shortcomings at the corresponding key exchange.
@@
Active Roaming Experiments. Large-scale studies that involve many operators and countries are usually limited by the complex ecosystem and the required coordination effort. However, there are several approaches (Mandalari2018Experience; Varvello2023Worldwide) where specifically built measurement devices were placed into target locations to measure the implications (e.g., QoE) of roaming. Additionally, Sahin and Francillon (Sahin2016Over) observed hijacking of traditional voice calls that were redirected to over-the-top (OTT) services (e.g., WhatsApp, Viber) to bypass/monetize termination fees.
@@
Recently, Gegenhuber et al. (Gegenhuber2023MobileAtlas; Gegenhuber2022Zero) introduced the MobileAtlas measurement platform that tries to overcome the mentioned scalability issues by tunneling the communication between SIM card and modem over the Internet. Their platform provides flexible roaming measurements and capabilities for a rich set of cellular features, including Internet and voice-based measurements.
@@
Evaluating and Fingerprinting VPNs. The commercial VPN ecosystem is a multi-billion dollar industry (VPNDemand) and thereby has been an interesting research target. For example, previous work (Khan2018Empirical; Ramesh2022Vpnalyzer) analyzed and compared existing VPN services and found that many solutions leak user traffic or advertise wrong server locations.
@@
In contrast, Maghsoudlou et al. (Maghsoudlou2023Characterizing) did not purchase any VPN subscriptions but executed Internet-wide scans to discover and fingerprint VPNs, finding over 7 million IPsec servers.
@@ 6. Discussion
Implications to Emergency Calling. Recent reports show, that there currently is no adequate support for VoLTE roaming in substantial parts of the world (Capacitymedia2022VoLTE1; Capacitymedia2022VoLTE2). Since many operators are actively shutting down their 2G/3G legacy networks, this scenario has the potential to result in significant repercussions for the functionality of emergency calling (Hendrik2022VoLTE; EENA2022VoLTE; EENA2022Webinar).
@@
Although we believe that affected operators will address and fix these issues in the long term, VoWiFi could help to mitigate these shortcomings in the present day. Tourists or travellers frequently have access to WiFi, such as in their accommodations or through free WiFi hotspots in public spaces. According to the specification (EtsiArchitectureNon3gpp), VoWiFi should be used by the UE for emergency calling when traditional radio services (VoLTE roaming, CSFB) are unavailable. The phone ultimately tries to reach both the home and the visited ePDG. However, there are circumstances where the visited ePDG might not be available, e.g., when there is no VoWiFi support in the visited country or when the customer leaves the phone in flight mode to not cause any unintentional roaming fees and thereby the current location is not known to the phone).
@@
While operators can override the default ePDG for emergency services by setting corresponding DNS records (sos.epdg.epc.mnc{y}.mcc{x}.pub.3gppnetwork.org(EtsiNumberingAdressingIdentification), only four operators had appropriate DNS entries. In all four cases, the emergency ePDG referenced the original ePDG’s IP address, which is also the default behaviour when no sos entry is found. Controversially, two of the operators specifically using sos-domains nevertheless block IKE-inquiries coming from foreign IP addresses.
@@
If an operator deploys DNS- or IKE-based geoblocking, these measures will also impact emergency services, actively denying persons in need from making an emergency call. Similarly, the emergency service via Wi-Fi is also unavailable in the home country, when customers use a VPN connection or an international SIM card (e.g., utilizing a travel router) that provides the Internet uplink via a foreign country.
Economic and Net Neutrality Perspective. In contrast to regular roaming over the radio interface – where the foreign roaming partner charges the home operator for the terminated calls and services – there is no additional economic overhead for VoWiFi calls that are initiated from overseas customers. In fact, calls terminated via VoWiFi are notably cost-effective for operators, as they eventually reduce expenses associated with the required radio transmission infrastructure (i.e., base stations) and spectrum licensing fees. Instead, the traffic is routed via external infrastructure (i.e., a WiFi AP) that was provided and paid for by the customer.
Moreover, adhering to the principle of net neutrality and the Open Internet, it is not allowed to discriminate (i.e., block) a customer’s data packets by their source or destination IP address, particularly when motivated solely by economic interests (cf. differential pricing and zero-rating). Assuming the discovered blocking practices were employed mainly for economic reasons, they could potentially be seen as a net neutrality violation.
Ways to Evade Geoblocking. Common mobile operating systems (i.e., Android and iOS) support changing the used DNS server for WiFi interfaces out of the box, which should allow easy bypassing of DNS-based blocking.
Additionally, both Android and iOS have built-in support for applying a system-wide VPN connection. However, even with an active VPN, VoWiFi uses a direct route over the WiFi interface. While this hinders standalone solutions on non-rooted phones, a viable alternative could be to use so-called travel routers. These devices act as an intermediary between WiFi AP and smartphone and typically offer the ability to redirect all traffic through a VPN connection, and thus via a customer’s home country.
Inaccurate Geolocation and Blocked VPN Services. Each VPN service uses one or more IP addresses for each country within its coverage. In the course of this study, we occasionally experienced outliers where certain IP addresses, address ranges, or Autonomous Systems (ASes) from specific VPNs experienced blocking even though belonging to the same country as the probed operator. We suppose that those connections were either blocked because their geolocation was wrongfully classified from the operator, or because the operator specifically blocked connections from IPs that are associated with certain VPN services.
@@ 6.1. Limitations
Although we are able to detect a considerable amount of blocking operators, several factors limit our approach. Due to the fact that these constraints limit the geoblocking variants that we’re able to detect, our results can be seen as a lower bound. Thereby, the actual occurrence of geoblocking in practice is most likely even higher.
@@
More Advanced Blocking. Our current approach is designed to detect simple IP-based blocking rules. However, some operators use more sophisticated (e.g., IMEI-based) ways to decide whether a customer should be able to use VoWiFi under current circumstances. For example, an operator could use the last known roaming status via the radio network to decide whether a customer is currently in their home country or abroad. Additionally, VoWiFi roaming can only be offered to a limited set of customers (i.e., premium users). For example, we found an Austrian operator that requires an additional subscription package to get VoWiFi enabled during international roaming (Magenta2023VoWiFiRoaming). In this case, the operator initially accepts IKE packets from all locations and decides whether to grant (or drop) access to VoWiFi at a later stage, when the subscriber’s identity is known. We focus on straightforward IP-based blocking because measuring these more advanced blocking techniques on a global scale would necessitate the acquisition of SIM cards for an unfeasible number of operators. Moreover, it is crucial to highlight that relying solely on IP address-based blocking for VoWiFi access also results in the restriction of access to emergency calling services. Implementing blocking techniques that compromise the unconditional availability of emergency services that people depend on in life-threatening situations is highly ill-advised.
@@
Implementation Incompatibility and Blocked VPN Services. Our IKE probing is based on a cellular-specific open-source implementation of the IKEv2 protocol [^10] https://github.com/fasferraz/SWu-IKEv2. Possibly, some ePDGs are not compatible with this implementation or do not answer requests that offer inappropriate cipher suites. Similarly, operators could block access from well-known VPN services or ASes and IP ranges that are used by them. As a countermeasure, we use multiple services to increase the diversity within our clients. Since we got responses from a high share of ePDGs that were discovered via DNS, these two factors do not considerably influence our results.
@@
Non-Native and Non-3GPP-Compliant VoWiFi. Whereas our approach focuses on the native 3GPP version of VoWiFi there are also other ways to support voice calling functionalities over Wi-Fi networks. For example, an operator might require their customers to download and install a dedicated VoWiFi app that gives direct access to the IMS network in a non-compliant way (GSMA2017ChinaVoWiFi). Additionally, some operators might use non-default ways to communicate their ePDGs, e.g., only resolve the hostname via an internal DNS server that is shipped to their customers via DHCP. Lastly, we do not cover VoIP calling via OTT services (e.g., WhatsApp, Viber, Skype).